# Arena: A Patch-of-Interest ViT Inference Acceleration System for Edge-Assisted Video Analytics

### v1

@@ 3.2. Keyframe Inference
(5) $\mathcal{M}_{a}\leftarrow\hat{\tilde{z}}_{0}^{1}=[\hat{x}^{1}_{p,1}\mathbf{E};\hat{x}^{1}_{p,2}\mathbf{E};\ldots;\hat{x}^{1}_{p,N}\mathbf{E}],$
@@
(6) $\hat{z}_{0}^{1}=\hat{\tilde{z}}_{0}^{1}+\mathbf{E}_{pos},\qquad\mathbf{E}_{pos}\in\mathbb{R}^{N\times D}.$
@@
Then, all tokens are processed through the $L$ layers transformer encoder to get the final output $\hat{z}_{L}^{1}$, which will be cached in the second memory token pool $\mathcal{M}_{b}$. The formation is:
(7) $\hat{z}_{\ell}^{1\prime}$ $=\operatorname{MSA}\left(\operatorname{LN}(\hat{z}^{1}_{\ell-1})\right)+\hat{z}^{1}_{\ell-1},$ $\ell=1\ldots L,$
$\hat{z}_{\ell}$ $=\operatorname{MLP}\left(\operatorname{LN}(\hat{z}_{\ell}^{1\prime})\right)+\hat{z}_{\ell}^{1\prime},$ $\ell=1\ldots L,$
(8) $\mathcal{M}_{b}\leftarrow\hat{z}_{L}.$
@@
Notably, memory token pools $\mathcal{M}_{a}$ and $\mathcal{M}_{b}$ will be updated in every frame inference and re-initialized at the next keyframe inference phase. Since the tokens of the keyframe are complete, we directly feed $\hat{z}_{L}$ into the MFR module. MFR is a single-layer transformer decoder (Vaswani et al., 2017) to build the final tokens and reshape to the strongest feature map $f_{3}$.
@@
We build a multi-scale feature pyramid $\{f_{1},f_{2},f_{3},f_{4}\}$ directly from intermediate tokens. $f_{3}$ and $f_{4}$ contain low-level information, while $f_{1}$ and $f_{2}$, after processing through the transformer encoder, contain high-level semantic information. Therefore, we combine these two sets of feature maps to enhance the effectiveness of object detection. Specifically, we use the token sequence before encoder (i.e., $\hat{\tilde{z}}_{0}^{1}$) and after MFR (i.e., $f_{3}$) as features, which are both at $1/16$ scale relative to the original frame. For $\hat{\tilde{z}}_{0}^{1}$, we use two sets of de-convolution operations to obtain feature maps $f_{1}$, $f_{2}$ at $1/4$ and $1/8$ scale. For $f_{3}$, we use the convolution of stride 2 to generate a feature map $f_{4}$ at $1/32$ scale. The formulations are
(9) $f_{1}=\text{DeConV}_{1}(\hat{\tilde{z}}_{0}^{1}),\qquad f_{2}=\text{DeConV}_{2}(\hat{\tilde{z}}_{0}^{1}),$
(10) $f_{3}=\text{MRF}(\hat{z}^{1}_{L}),\qquad f_{4}=\text{ConV}(f_{3}).$
@@
Finally, $\{f_{1},f_{2},f_{3},f_{4}\}$ from the ViT backbone form a multi-scale feature pyramid, which is then fed into two-stage detectors, such as Faster R-CNN (Ren et al., 2015) or Cascade R-CNN (Cai and Vasconcelos, 2019), to generate the bounding boxes and categories of the objects.
3.3. Non-keyframe Inference with Memory Feature Reconstruction
@@
During non-keyframe inference, the PPS first identifies $N^{\prime}$ PoIs using the last frame detection results (e.g., results from $\hat{\mathbf{x}}^{1})$. Then, the camera only extracts the PoIs out of the frame and offloads them to the edge server, which is denoted as $\mathbf{x}_{p}^{2}=\{x^{2}_{p,1},\ldots,x^{2}_{p,N^{\prime}}|x^{2}_{p,i}\in\mathbb{R}^{P^{2}C}\}$. Here $N^{\prime}\ll N$ due to the fact that objects only occupy a very small portion of the frame. After linear projection for $\mathbf{x}_{p}^{2}$, we get a sparse token sequence $\tilde{z}_{0,sp}^{2}$. Subsequently, we add position embeddings corresponding to the original frame position to each token as
(11) $z_{0,sp}^{2}=\tilde{z}_{0,sp}^{2}+\mathbf{E}^{\prime}_{pos,sp},\qquad\mathbf{E}^{\prime}_{pos,sp}\in\mathbb{R}^{N^{\prime}\times D}.$
Next, $z_{0,sp}^{2}$ is fed into the transformer encoder to compute $z_{L,sp}^{2}$ following Eq. (7).
@@
To restore the complete token sequence from the sparse one, we need to reuse the tokens in memory token pools. We extract tokens from non-PoI regions in $\mathcal{M}_{a}$ and combine them with the PoIs token $\tilde{z}_{0,sp}^{2}$ to restore a complete token sequence $\tilde{z}_{0}^{2}$. Similarly, we extract tokens from non-PoI regions in $\mathcal{M}_{b}$ and combine them with $z_{L,sp}^{2}$ to restore $z_{L}^{2}$. For example, as illustrated in Figure 3, the edge server only receives and processes tokens at positions 1, 2, 3, 5, and 6. Then, we insert the tokens from the remaining positions of $\mathcal{M}_{a}$ and $\mathcal{M}_{b}$ (i.e., 4, 7, 8, and 9) into the processed tokens $\tilde{z}_{0,sp}^{2}$ and $z_{L,sp}^{2}$ to restore complete sequences, respectively. Meanwhile, we introduce MFR to enhance the PoI tokens with background information from historical frames with minimal computational overhead. Finally, we follow the Eq. (9) and Eq. (10) to get the multi-scale feature pyramid $\{f_{1},f_{2},f_{3},f_{4}\}$.
@@
(12) $\mathcal{M}_{a}\leftarrow\tilde{z}_{0}^{2},\mathcal{M}_{b}\leftarrow z_{L}^{2}.$
3.4. Probability-based Patch Sampling
@@
Figure 4. Our probability-based patch sampling mechanism. The blue patch indicates that it is highly likely to be determined as an PoI. Zoom in for the best view.
@@
How to determine the PoIs in the next frame is a critical challenge. Too few patches may cause the model to miss detections, while too many patches could transmit unnecessary background information and add extra computational overhead. To address this issue, we design a lightweight and effective probability-based patch sampling mechanism that leverages history information to determine the PoIs. Our insight is that the region where the object appears in the next frame will be near the bounding box of the previous frame with high probability.
@@
As shown in Figure 4, the two frames are adjacent. In the first frame $\mathbf{x}^{k}$, we have obtained the bounding box of the objects (box in red), and all patches occupied by this bounding box (patches in red) are considered as the PoIs of the previous frame. Next, we expand the PoIs region by appending $m$ (referred to as the expanded bbox margin) rows and columns around the perimeter, including the top, bottom, left, and right edges (patches in yellow). These patches are regarded as the sampling region $R_{s}$ for the current frame $\mathbf{x}^{k+1}$ (patches in blue). We use inter-frame differences to determine the regions that need additional transmission to alleviate the miss of new objects that enter the scene for the first time. Specifically, we convert $\mathbf{x}^{k}$ and $\mathbf{x}^{k+1}$ to greyscale images and compute the pixel difference $f_{d}$ between the two greyscale images. Next, we divide the greyscale image $f_{d}$ into non-overlapping $16\times 16$ patches, identifying patches outside the $R_{s}$ with pixel differences exceeding the threshold $F$. The new objects are likely to be found in these patches, which are also marked as sampling regions. Finally, we randomly sample $p\%$ of patches from the sampling region as PoIs because our backbone is not sensitive to whether the object area is complete, and the background can be reconstructed using MFR.
3.5. Training Procedure
@@
We fine-tune the whole detector based on a pre-trained ViT backbone in a video-training manner. During fine-tuning, for each training batch, we input a pair of adjacent frames $\{\hat{\mathbf{x}}^{k},\mathbf{x}^{k+1}\}$ to the model to simulate our keyframe and non-keyframe inference procedure. Both frames use identical preprocess pipeline, including data augmentation techniques such as cropping and flipping. After that, we take the first frame $\hat{\mathbf{x}}^{k}$ as the keyframe and feed all patches from $\hat{\mathbf{x}}^{k}$ into the ViT backbone to initialize the memory token pools $\mathcal{M}_{a}$ and $\mathcal{M}_{b}$. Since directly using predicted bounding boxes as results may disturb the proper detection in the next frame during training, we use the ground truth bounding boxes instead of predicted ones from the first frame as the detection results. Finally, for the second frame $\mathbf{x}^{k+1}$, we follow Sec. 3.4 to get the PoIs and Sec. 3.3 to conduct a non-keyframe inference. It’s worth noting that we only use the inference results of the second frame for back-propagation and network updates. In addition, we utilize the pre-trained weights of ViT backbone derived from masked image modeling techniques (He et al., 2022; Ren et al., 2023).
4. Implementation

### v2

@@ III-B Keyframe Inference
$\mathcal{M}_{a}\leftarrow\hat{\tilde{z}}_{0}^{1}=[\hat{x}^{1}_{p,1}\mathbf{E};\hat{x}^{1}_{p,2}\mathbf{E};\ldots;\hat{x}^{1}_{p,N}\mathbf{E}],$ (5)
@@
$\hat{z}_{0}^{1}=\hat{\tilde{z}}_{0}^{1}+\mathbf{E}_{pos},\qquad\mathbf{E}_{pos}\in\mathbb{R}^{N\times D}.$ (6)
@@
Then, all tokens are processed through $L$ transformer blocks to get the final output $\hat{z}_{L}^{1}$, which will be cached in the second memory token pool $\mathcal{M}_{b}$. The formation is
$\hat{z}_{\ell}^{1\prime}$ $=\operatorname{MSA}\left(\operatorname{LN}(\hat{z}^{1}_{\ell-1})\right)+\hat{z}^{1}_{\ell-1},$ $\ell=1\ldots L,$ (7)
$\hat{z}_{\ell}^{1}$ $=\operatorname{MLP}\left(\operatorname{LN}(\hat{z}_{\ell}^{1\prime})\right)+\hat{z}_{\ell}^{1\prime},$ $\ell=1\ldots L,$
$\mathcal{M}_{b}\leftarrow\hat{z}_{L}^{1}.$ (8)
@@
Notably, memory token pools $\mathcal{M}_{a}$ and $\mathcal{M}_{b}$ will be updated in every frame inference and re-initialized at the next keyframe inference phase. Since the tokens of the keyframe are complete, we directly feed $\hat{z}_{L}^{1}$ into the MFR module. MFR is a single-layer transformer decoder [31] to build the final tokens and reshape to the strongest feature map $f_{3}$.
@@
We build a multi-scale feature pyramid $\{f_{1},f_{2},f_{3},f_{4}\}$ directly from intermediate tokens. $f_{1}$ and $f_{2}$ contain low-level information, while $f_{3}$ and $f_{4}$, after processing through the transformer encoder, contain high-level semantic information. Therefore, we combine these two sets of feature maps to enhance the effectiveness of object detection. Specifically, we use the token sequence before encoder (i.e., $\hat{\tilde{z}}_{0}^{1}$) and after MFR (i.e., $f_{3}$) as features, which are both at $1/16$ scale relative to the original frame. For $\hat{\tilde{z}}_{0}^{1}$, we use two sets of de-convolution operations to obtain feature maps $f_{1}$, $f_{2}$ at $1/4$ and $1/8$ scale. For $f_{3}$, we use the convolution of stride 2 to generate a feature map $f_{4}$ at $1/32$ scale. The formulations are
$f_{1}=\text{DeConV}_{1}(\hat{\tilde{z}}_{0}^{1}),\qquad f_{2}=\text{DeConV}_{2}(\hat{\tilde{z}}_{0}^{1}),$ (9)
$f_{3}=\text{MFR}(\hat{z}^{1}_{L}),\qquad f_{4}=\text{ConV}(f_{3}).$ (10)
@@
Finally, $\{f_{1},f_{2},f_{3},f_{4}\}$ from the ViT backbone form a multi-scale feature pyramid, which is then fed into two-stage detectors, such as Faster R-CNN [32] or Cascade R-CNN [33], to generate the bounding boxes and categories of the objects.
III-C Non-keyframe Inference with Memory Feature Reconstruction
@@
During non-keyframe inference, the PPS first identifies $N^{\prime}$ PoIs using the last frame detection results (e.g., results from $\hat{\mathbf{x}}^{1})$. Then, the camera only extracts the PoIs out of the frame and offloads them to the edge server, which is denoted as $\mathbf{x}_{p}^{k}=\{x^{k}_{p,1},\ldots,x^{k}_{p,N^{\prime}}|x^{k}_{p,i}\in\mathbb{R}^{P^{2}C},k\in(1,K],i\in[1,N^{\prime}]\}$. Here $N^{\prime}\ll N$ due to the fact that objects only occupy a very small portion of the frame. After linear projection for $\mathbf{x}_{p}^{k}$, we get a sparse token sequence $\tilde{z}_{0,sp}^{k}$. Subsequently, we add position embeddings corresponding to the original frame position to each token as
$z_{0,sp}^{k}=\tilde{z}_{0,sp}^{k}+\mathbf{E}^{\prime}_{pos,sp},\qquad\mathbf{E}^{\prime}_{pos,sp}\in\mathbb{R}^{N^{\prime}\times D}.$ (11)
Next, $z_{0,sp}^{k}$ is fed into the transformer encoder to compute $z_{L,sp}^{k}$ following Eq. (7).
@@
To restore the complete token sequence from the sparse one, we need to reuse the tokens in memory token pools. We extract tokens from non-PoI regions in $\mathcal{M}_{a}$ and combine them with the PoIs token $\tilde{z}_{0,sp}^{k}$ to restore a complete token sequence $\tilde{z}_{0}^{k}$. Similarly, we extract tokens from non-PoI regions in $\mathcal{M}_{b}$ and combine them with $z_{L,sp}^{k}$ to restore $z_{L}^{k}$. For example, as illustrated in Figure 5, the edge server only receives and processes tokens at positions 1, 2, 3, 5, and 6. Then, we insert the tokens from the remaining positions of $\mathcal{M}_{a}$ and $\mathcal{M}_{b}$ (i.e., 4, 7, 8, and 9) into the processed tokens $\tilde{z}_{0,sp}^{k}$ and $z_{L,sp}^{k}$ to restore complete sequences, respectively. Meanwhile, we introduce MFR to enhance the PoI tokens with background information from historical frames with minimal computational overhead. Finally, we follow the Eq. (9) and Eq. (10) to get the multi-scale feature pyramid $\{f_{1},f_{2},f_{3},f_{4}\}$.
@@
$\mathcal{M}_{a}\leftarrow\tilde{z}_{0}^{k},\mathcal{M}_{b}\leftarrow z_{L}^{k}.$ (12)
III-D Probability-based Patch Sampling
@@
Figure 6: Our probability-based patch sampling mechanism. The blue patch indicates that it is highly likely to be determined as an PoI. Zoom in for the best view.
Input: Previous frame $\mathbf{x}^{k}$, current frame $\mathbf{x}^{k+1}$, pixel diffrences thereshold $F$, sampling rate $p$, expanded bbox margin $m$, bounding boxes $B^{k}=\{[\mu_{1,1}^{k},\nu_{1,1}^{k},\mu_{2,1}^{k},\nu_{2,1}^{k}],...\}$ in previous frame.
Output: Sampled patches ${R}^{k}_{s}$
1 $R^{k}_{s}\leftarrow[\emptyset]$ initialize the sampled patch list;
2 for $b\in B^{k}$ do
3       $P^{k}\leftarrow\textit{bbox\_to\_PoI}(b)$;
4       $P^{k}_{s}\leftarrow\textit{expand\_PoI}(P^{k},m)$;
5       $R^{k}_{s}.\textit{append}(P^{k}_{s})$;
6
7 end for
[…]
9 for $f^{k}_{i}\in\textbf{f}^{k}$ do
10       if $\textit{sum}(f^{k}_{i})>F\land f^{k}_{i}\notin R^{k}_{s}$ then
11             $R^{k}_{s}.\textit{append}(f^{k}_{i})$;
12
13       end if
14
15 end for
[…]
17 Function $(b)$:
18       $\mu_{1}^{k}\leftarrow{\mu_{1}^{k}//16\times{16}}$;
19       $\nu_{1}^{k}\leftarrow{\nu_{1}^{k}//16\times{16}}$;
20       $\mu_{2}^{k}\leftarrow{(\mu_{2}^{k}//16+1)\times{16}}$;
21       $\nu_{2}^{k}\leftarrow{(\nu_{2}^{k}//16+1)\times{16}}$;
22
23 return $[\hat{x}_{1}^{k},\hat{y}_{1}^{k},\hat{x}_{2}^{k},\hat{y}_{2}^{k}]$
24 Function expand_PoI$(P^{k},m)$:
25       $\widetilde{\mu}_{1}^{k}\leftarrow{\hat{\mu}_{1}^{k}-m\times 16}$;
26       $\widetilde{\nu}_{1}^{k}\leftarrow{\hat{\nu}_{1}^{k}-m\times 16}$;
27       $\widetilde{\mu}_{2}^{k}\leftarrow{\hat{\mu}_{1}^{k}+m\times 16}$;
28       $\widetilde{\nu}_{2}^{k}\leftarrow{\hat{\nu}_{1}^{k}+m\times 16}$;
29
30 return $[\widetilde{\mu}_{1}^{k},\widetilde{\nu}_{1}^{k},\widetilde{\mu}_{2}^{k},\widetilde{\nu}_{2}^{k}]$
31
Algorithm 1 Probability-based Patch Sampling
@@ 28       $\widetilde{\nu}_{2}^{k}\leftarrow{\hat{\nu}_{1}^{k}+m\times 16}$;
Determining the PoIs in the next frame is a critical challenge. Too few patches may cause the model to miss detections, while too many patches could transmit unnecessary background information and add extra computational overhead. To address this issue, we design a lightweight and effective probability-based patch sampling mechanism that leverages history information to determine the PoIs, shown in Algorithm. 1. Our insight is that the region where the object appears in the next frame will be near the bounding box of the previous frame with high probability.
@@
As shown in Figure 6, the two frames are adjacent. zIn the first frame $\mathbf{x}^{k}$, we have obtained the bounding box of the objects (box in red), and all patches occupied by this bounding box (patches in red) are considered as the PoIs of the previous frame. Next, we expand the PoIs region by appending $m$ (referred to as the expanded bbox margin) rows and columns around the perimeter (Lines 1-6), including the top, bottom, left, and right edges (patches in yellow). These patches are regarded as the sampling region $R_{s}^{k}$ for the current frame $\mathbf{x}^{k+1}$ (patches in blue). We use inter-frame differences to determine the regions that need additional transmission to alleviate the miss of new objects that enter the scene for the first time. Specifically, we convert $\mathbf{x}^{k}$ and $\mathbf{x}^{k+1}$ to greyscale images and compute the pixel difference $\mathbf{f}^{k}$ between the two greyscale images (Line 7). Next, we divide the greyscale difference image $\mathbf{f}^{k}$ into non-overlapping $16\times 16$ patches ($\mathbf{f}^{k}=\{{f}^{k}_{1},\cdots,f^{k}_{N}\}$) identifying patches outside the $R_{s}^{k}$ with pixel differences exceeding the threshold $F$ (Lines 8-12). The new objects are likely to be found in these patches, which are also marked as sampling regions. Finally, we randomly sample $p\%$ of patches from the sampling region as PoIs (Line 13) because our backbone is not sensitive to whether the object area is complete, and the background can be reconstructed using MFR.
III-E Adaptive Keyframe Interval Switching Algorithms
Input: Optical flow threshold $\beta$, keyframe $\hat{\mathbf{x}}_{r}^{1}$, last frame in the interval $\mathbf{x}_{r}^{K}$, current keyframe interval $K_{r}$, bounding boxes $B_{r}^{1}$ in keyframe.
Output: Next keyframe interval $K_{r+1}$
1 $V_{bbox}\leftarrow 0,A_{bbox}\leftarrow 0$;
2 $\mathbf{V}\leftarrow\textit{caculate\_optical\_flow\_diff}(\hat{\mathbf{x}}_{r}^{1},\mathbf{x}_{r}^{K})$;
3 $R_{b}\leftarrow\textit{zeroes}(\mathbf{V}.\textit{row},\mathbf{V}.\textit{column})$ initialize a mask matrix of the same shape as $\mathbf{V}$;
4 for $b\in B_{r}^{1}$ do
5       $R_{b}(b)\leftarrow 1$;
6
7 end for
[…]
9 $A_{bbox}\leftarrow Area(R_{b})$;
10 $\overline{V}_{bbox}\leftarrow\frac{V_{bbox}}{A_{bbox}}$;
11 if $\overline{V}_{bbox}<\beta\land K_{r}<K_{upper}$  then
12       $K_{r+1}\leftarrow K_{r}+1$
13else if $\overline{V}_{bbox}>\beta\land K_{r}>K_{lower}$  then
14       $K_{r+1}\leftarrow K_{r}-1$
15 end if
16return $K_{r+1}$;
Algorithm 2 Adaptive Keyframe Interval Switching
Keyframe interval introduces a trade-off between accuracy and bandwidth, as previously discussed in Sec. II-E. A promising method for adjusting keyframe interval is to assign a smaller value to those “complex” videos with rapid changes (e.g., dash camera) and a larger value to those “simple” videos with slower changes (e.g., surveillance camera). To do this, we propose the AKIS algorithms to control the keyframe interval to achieve optimal accuracy and bandwidth trade-off. Inspired by the optical flow method [34], our insight is to set a threshold for optical flow difference values, using it as a bidirectional switch to adjust the keyframe interval.
The AKIS algorithm is designed to determine the length of the r+1-th keyframe interval $K_{r+1}$ based on the optical flow information from the current interval $K_{r}$, shown in Algorithm. 2. Specifically, we calculate the optical flow map $\mathbf{V}$ between the keyframe $\hat{\mathbf{x}}_{r}^{1}$ and the last non-keyframe $\mathbf{x}_{r}^{K}$ in the current interval (Line 2). The magnitude of an optical flow map represents the extent of pixel motion between two frames. To mitigate the impact of noise (i.e., background pixel, camera shifts) on the optical flow map, we retain only the non-overlapping bounding box regions in keyframes to calculate the magnitude $\mathbf{V}(b)$ (Lines 3-8). Next, the average optical flow magnitude $\overline{V}_{bbox}$ is calculated as the total magnitude divided by the non-overlapping area of all bounding boxes (Line 9). Ultimately, a comparison is made between $\overline{V}_{bbox}$ and the threshold $\beta$. If $\overline{V}_{bbox}$ is lower than $\beta$, the object changes in the current frame interval are slow, which can be considered a “simple” scene. Consequently, the length of the subsequent interval $K_{r+1}$ is increased by one (Lines 10-11). On the contrary, if the “complex” scene is determined, the length of the next interval $K_{r+1}$ is decreased by one (Lines 12-13).
The algorithm will return a new keyframe interval, thereby enabling the transmission bandwidth to be modified in real time for different scenes. In scenes where the rate of change for objects or people is minimal, the algorithm can extend the keyframe interval, thereby reducing bandwidth usage between the camera and the edge server while maintaining inference accuracy. Conversely, in high-change-rate scenes, the algorithm can continuously decrease the keyframe interval to ensure precision. By leveraging the trade-off relationship, this algorithm enables an adaptive video inference system that dynamically balances accuracy and bandwidth.
III-F Training Procedure
@@
We fine-tune the whole detector based on a pre-trained ViT backbone in a video-training manner. During fine-tuning, for each training batch, we input a pair of adjacent frames $\{\hat{\mathbf{x}}^{k},\mathbf{x}^{k+1}\}$ to the model to simulate our keyframe and non-keyframe inference procedure. Both frames use identical preprocess pipeline, including data augmentation techniques such as cropping and flipping. After that, we take the first frame $\hat{\mathbf{x}}^{k}$ as the keyframe and feed all patches from $\hat{\mathbf{x}}^{k}$ into the ViT backbone to initialize the memory token pools $\mathcal{M}_{a}$ and $\mathcal{M}_{b}$. Since directly using predicted bounding boxes as results may disturb the proper detection in the next frame during training, we use the ground truth bounding boxes instead of predicted ones from the first frame as the detection results. Finally, for the second frame $\mathbf{x}^{k+1}$, we follow Sec. III-D to get the PoIs and Sec. III-C to conduct a non-keyframe inference. It’s worth noting that we only use the inference results of the second frame for back-propagation and network updates. In addition, we utilize the pre-trained weights of ViT backbone derived from masked image modeling techniques [18].
IV Implementation